Sol Summary – April 19th

Crew 210 Sol Summary Report 19-APR-2019

Sol: 6

Summary Title: Writing and preparing for crew handover

Author’s name: Paul Sokoloff

Mission Status: Nominal

Sol Activity Summary: Today the crew prepared to depart MDRS and leave it to our successors on Crew 211. We spent the morning cleaning the station, inventorying the food, and getting the spaces ship-shape for handover.

In the afternoon we continued to process plant samples and salinity measurements in the Science Dome, and collaboratively wrote a white paper outlining our multi-year biodiversity program, while Mike Irvine prepared an in-depth outreach proposal.

We will continue to prepare for our imminent departure this evening.

Look Ahead Plan: We plan to leave MDRS at 0800 tomorrow to head to Grand Junction, where we plan on handing over the crew car to Crew 211 by 1200.

Thank you to Mission Support, the Mars Society, and the management of MDRS for a great week in Utah; the five of us really appreciate your support.

Anomalies in work: No problems.

Weather: Hot and sunny.

Crew Physical Status: Good.

EVA: No EVAs today.

Reports to be filed: GreenHab Report, Operations Report, Daily Photo Report, Mission Summary, End of Rotation Food Inventory, End of Rotation Checklist (by tomorrow am).

Support Requested: None

Greenhab Report – April 19th

GreenHab Officer: David Murray

Environmental control: Ambient with fan 0930-1900

Shade Cloth: 80%

Low Temperature: 15 C

High Temperature: 37 C

Relative humidity: 15%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 34 L total, 8 L fertilizer

Water in Blue Tank: refilled with about 85 gallons of water

Time(s) of Watering: 1100 and 1850

Changes to Crops: Tomatoes are stressed by the heat (curling, drooping, yellowing of basal foliage)

Narrative: It was very hot out today and the tomatoes didn’t like that. The other plants look healthy. I fertilized again because not everything was fertilized the first time and the first application was light. The fan was on from 0930 to 1900 but was turned off while filming a GreenHab video.

Tonight was my last night as the GreenHab officer because Crew 210 departs for Grand Junction at 0800 tomorrow morning.

Harvest: None

Operations Report – April 19th

Crew 210 Operations Report 19-APR-2019

SOL: 7

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 12

On (last night): 7 pm

Off (this morning): 7 am

Solar— SOC 75 % at 7 pm (so turned generator on)

Diesel Reading – 34 %

Propane Reading – 49%

Ethanol Free Gasoline – 4 gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 425 gallons

Auxiliary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 150 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours: Did not record

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours: did not check

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: not recorded

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours: did not record

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) all

Reason for use: moving them back over to Shannon’s house

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? Yes, a trip to town

General notes and comments: Submersible pump in the transfer water tank is working

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Cleaning

Summary of GreenHab operations: David will send a full report today.

Summary of ScienceDome operations: Finishing up lab work and cleaning

Summary of RAM operations: Cleaning

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing tonight,
thank you.

Copyright © The Mars Society. All rights reserved. | Main Site